OpenStudy (anonymous):

Write the equation of the cosine function with amplitude 2 period 4p

OpenStudy (anonymous):

Amplitude* cos(2pi/period)

OpenStudy (anonymous):

period of \[cos(bx)\] is \[\frac{2\pi}{b}\] here you want period of \[4\pi\] so set \[\frac{2\pi}{b}=4\pi\] and solve for \[b\] to get \[b=\frac{1}{2}\]

OpenStudy (anonymous):

so your function will be \[2cos(\frac{x}{2})\]

OpenStudy (anonymous):

much faster this afternoon yes?

OpenStudy (amistre64):

"Write the equation of the cosine function with amplitude 2; period 4p" This depends on what exactly they mean by 4p, maybe (4pi?) but: 2 cos(x/2)

OpenStudy (amistre64):

little but, they must have replaced the Dcells in the ti-83 lol